OpenStudy (anonymous):

please help will give medal! what is the value of a? 9a-3(a-6)= -6

OpenStudy (anonymous):

First get rid of your parenthesis

OpenStudy (anonymous):

9a-3a-6= -6 ?

OpenStudy (anonymous):

You have to multiply everything in the parenthesis by -3, so try again

OpenStudy (anonymous):

i dont get it

OpenStudy (anonymous):

I'll show you

OpenStudy (anonymous):

You have to distribute the -3 to the insides of the parenthesis by multiplication for them to be taken away. your equation is:\[9a-3(a-6)=-6 \] so by distributing and multiplying everything in the parenthesis, you will end up with: \[9a-3a+18=-6\]

OpenStudy (anonymous):

ohh okay

OpenStudy (anonymous):

Now subtract 3a from 9a, then after you do that subtract 18 from both sides, then divide both sides by 6

OpenStudy (anonymous):

so after the subtracting would it be 6a+18= -6?

OpenStudy (anonymous):

@megan_tanuis

OpenStudy (anonymous):

Yes, that's correct. Now subtract 18 from both sides